OpenStudy (anonymous):

Simplify the following expression: 3(3z - 4 ) - (20 z - 5) Answer: a. 11 z - 7 b. 11z + 17 c. - 11z - 7 d. - 11z - 17 e. - 51z + 3

OpenStudy (anonymous):

sattelite will you help me please

OpenStudy (anonymous):

You first have to distribute the 3. 3(3z - 4 ) - (20 z - 5) And distribute the negative sign also 9z - 12 - 20z + 5 Then combine like terms.

OpenStudy (anonymous):

9z-12-20z+5=0 -11z-7=0 11z+7=0

OpenStudy (anonymous):

9z - 20z and -12 + 5 -11z - 7

OpenStudy (anonymous):

A rectangular display case houses a square pyramid. Both have the same square base measuring x inches on a side. The top of the case is 2 inches above the tip of the pyramid which has a height of h inches. Which expression represents the volume of the empty space inside the case? a. x^3 - x^2(h-2) b. x^2-1/3x^2h c. 1/3x2^h d. 2/3x^2h e. x^2(h+2) - 1/3x^2h

OpenStudy (anonymous):

will someone help please

OpenStudy (anonymous):

its c...

OpenStudy (anonymous):

yup he is right ... its C try it ur self and see : ]

OpenStudy (anonymous):

actually it ll form a cone..n u hve to fnd its vol.

OpenStudy (anonymous):

If Rob flips 3 coins what is the chance that they will come up all tails? Answer: a. 1/2 b. 1/4 c. 1/6 d. 1/8 e. 1/16

OpenStudy (anonymous):

this is my last question will someone help me